OpenStudy (anonymous):

isaiah sells tennis equipment. he is guaranteed a minimum salary of $1500 per month plus 5.75% off his total sales. what are his total sales for a month in which his gross pay is $2075

OpenStudy (anonymous):

anyone know??

OpenStudy (anonymous):

Gross pay-min saqlary = 2075 -1500 = 575. this figure($575) must represent the 5.75% he got from his total sales. so, if 5.75% = $575 then to calculate 1%, we divide both sides by 5.75, to give us 1%=575/5.75 = $100 therefore, 100% of his sales is =$10,000 so to check we can say $10,000/100 = 1% i.e. 1% = 100 and 5.75% of this is 5.75 x$100 = $575